# Quantum determinism and completeness restored by indistinguishability and long-time particle detection

### v1

@@ IV.3.1 Superconducting qubit inside a cavity
Quantum non-demolition (QND) measurement is one of the current techniques for monitoring few-level systems (qubit or cavity photon). In a quantum circuit, a qubit readout is achieved by an interaction with a detuned cavity allowing a quantum non-demolition measurement in the dispersive regime. If a qubit is in an initial state $|\psi\rangle_{1}=c_{0}|0\rangle+c_{1}|1\rangle_{V}$, an external input radiation allows, after interaction, the transmission of an ideal coherent state $|\alpha\rangle$ inside a cavity mode for the excited state, the final “cat-like” state $|\psi\rangle_{1}=c_{0}|0\rangle|0\rangle+c_{1}|1\rangle|\alpha\rangle$ [25, 26]. The resulting output signal has two possible unambiguous intensity outcomes 00 or $I=|\alpha|^{2}$ in absence of a superposition. Considering $N$ measurements without a detector reset in between two events, the average total intensity is $\overline{I}=N|c_{1}|^{2}I$ and the fluctuations $\overline{\delta^{2}I}=N[|c_{1}|^{2}(1-|c_{1}|^{2})]I^{2}$. Clearly, the relative fluctuations shrink to zero for large $N$. In this sense, $|c_{0}|^{2}$ and $|c_{1}|^{2}$ are the probability amplitude that is measurable after a counting over many events.
@@ V.1 General considerations
We neither claim that quantum correlations or entanglement at long distance are not possible, nor that Bell states do not exist. The singlet or triplet states in the Helium atoms are examples of such entangled electron spin pairs. However, performing any Bell-like experiment in a framework based on a many-mode description is questionable once it is established that the collected data correspond to events that are not well separated because of their space and time correlations. Hence, there cannot be independent measurements in the context of a Bell experiment. Such a type of experiment would certainly be feasible if there would be a means to distinguish identical particles, which would substantiate the assumption that two consecutive measurement events can be well separated. But given the experimental reality of non-discrimination between the detector and the setup, the current interpretation of Bell experiments does not involve the physical description of the detector and the source. Without the Born’s postulate, the quantum description of the Bell experiment needs be reformulated. Particularly, the impossibility of a perfect interpretation of a noisy measurement prevents its feasibility, and hence any of two non-local correlated observed outcomes (spin up, spin down) is necessarily the result of many non separable events. Under these conditions, locality, causality, and completeness are restored and there is no need of additional, hidden variables as shown below.
@@ V.2 Influence of the source thermal noise on photon pair production
A set of pair states produced in an experiment is quite different from what John Bell assumed in his work [27]. Using a continuous variable formalism associated to photons, we can factorize the polarized state symmetrically if the EPR source is not noisy. But in this case, all outcomes in both polarisation directions would occur simultaneously contrary to observations where only one direction is observed. A possible explanation for such a discrimination is the inherent thermal noise of the source, like a beam of calcium atoms [14, 15, 16] or a nonlinear crystal [18] that generates additional classical photon flux fluctuations blurring the intrinsic quantum ones, creating a set of random pulse events.
In a particle picture, an itinerant photon shows a Brownian-like dynamics as it interacts with atoms whose motion relative to the detector produces a Doppler shift with a relative frequency change $\sim\sqrt{1/\beta m}/c\sim 10^{-6}$ where $m$ is the mass of the atoms of the source, $c$ the speed of light, and $\beta$ the inverse thermal energy reflecting the temperature statistically determined from the average over the speeds of the atoms in motion.
@@
We can consider it as a stochastic variable whose relative average variation is given by the Doppler shift of a moving atom within the nonlinear crystal or the velocity distribution of calcium atoms. This results in a variance given by $\overline{\eta^{2}_{k,\epsilon,\epsilon^{\prime}}}(t)=\delta_{\epsilon,\epsilon^{\prime}}(\hbar\omega_{k})^{2}/\beta mc^{2}$ Comparing with the detuning and the nonlinear coefficient, these stochastic terms have an effect $(\omega_{k_{0}}/\Delta_{k_{0}})\sqrt{1/\beta mc^{2}}$, which can be of the order of unity or greater. As a consequence, such important contributions lead to potentially strong local variations in the intensity for creating histories of polarization mimicking hidden variables.
A basic formulation of the Bell experiment relies on the multi-mode squeezed state:
@@
Assuming instead that pulses with photon numbers greater than unity are negligible with only one wavevector mode, we could employ the naive approximation of projecting the state onto the particle number pair of Eq. (20) to recover the symmetric Bell entangled state: $|\psi\rangle=(\hat{a}^{\dagger}_{k,H}\hat{b}^{\dagger}_{-k,H}+\hat{a}^{\dagger}_{k,V}\hat{b}^{\dagger}_{-k,V})|0\rangle/\sqrt{2}$. The use of the Born’s probability postulate allows the possibility of different polarisation outcomes.
@@ VI Conclusion
We also discussed with basic illustrative examples the use of the second quantization multi-particle formalism to avoid the pitfalls of the wavepacket collapse in the Schrodinger’s cat paradox. While the use of statistics is inherent to quantum physics experiments, this does not necessary entail that probability is a basic feature of quantum mechanics. It is inherent because the intrinsic uncertainty associated to common observables including, e.g., position, momentum, or particle number, requires a large number of data to be properly defined as a quasi-noiseless macroscopic measurement, which is essential for a conclusive interpretation. This amounts to adopting a thermodynamics formulation of the problem, where only macroscopic quantities are measurable since any single-event signal is too noisy to ensure a valuable characterization. For example, the observation of the electronic cloud in a solid crystal under a scanning tunnelling microscope is not a reality proposed by quantum mechanics but a representation of the total electronic density distribution resulting from of a macroscopic measurement current.
From our proposed many-body modeling including the thermal noise of the source and of the detector, we conclude that the experiments in Refs. [14, 15, 16, 20, 18] are not proven either to correspond to the idealistic Bell’s proposal. Any genuine modeling of a quantum correlation experiment, with the hope of more deterministic predictions, should ideally include: 1) The full formalism of quantum electrodynamics with a many-body formalism highlighting indistinguishability; 2) A correct modelling of the source (e.g., a calcium beam with the two lasers); 3) An accurate quantum model of the detector which operates in a finite time and whose finite size imposes a lower bound over the counting rate preventing genuinely distinguishable (time-separated) single-event detection. A procedure along these steps results in a deterministic quantum description including locality and causality.
@@
More complex experiments that necessitate modeling beyond the traditional scattering approach using the S matrix, such as, e.g., the avalanche effect, need also new theoretical methods to address properly the exact quantum dynamics involved. For instance, the formalism of transmission within wave-guides usually derived from classical electromagnetism has been extended using quantization [21]. These developments raise the fundamental questions on the possibility of creating a superposition of macroscopic “cat” states leading to large quantum fluctuations and therefore not measurable. In this work, we have shown the difficulty to produce such a “multiverse” state given the presently available technology; but a more accurate proof is required to establish this observation on a more formal basis. The non existence of such states amounts to conjecturing that any macroscopic observable has its set of possible eigenvalues restricted to regions within its fluctuations, discarding any extreme value far from its average. Such a statement should ideally be proven starting from a quantum field model.
@@
As a final remark regarding the discrimination in the theoretical description between the object of study and the measurement apparatus, it is worth recalling that when electric circuits were discovered at the time of Alessandro Volta in 19th century, both the circuit (e.g. voltaic pile + metal wires) and the measurement device (e.g. compass or galvanometer) had eventually required the electromagnetic theory for a full understanding of an experiment.

### v2

@@ IV.3.1 Superconducting qubit inside a cavity
Quantum non-demolition (QND) measurement is one of the current techniques [58] for monitoring few-level systems (qubit or cavity photon). In a quantum circuit, a qubit readout is achieved by an interaction with a detuned cavity allowing a quantum non-demolition measurement in the dispersive regime. Concretely, the qubit readout involves the measurement of a macroscopic current through a microwave waveguide and uses the Stark shift effect to discriminate between one state or another in a specific basis [57]. But in the case of a superposition, there is no such detailed experimental analysis providing evidence that the Born rule applies. It may well be that the outcome is such that no discrimination is possible between one eigenstate over another (possibly leading to dark counting) and still applying Born’s rule [7] while its validity is questionable.
If a qubit is in an initial state $|\psi\rangle_{1}=c_{0}|0\rangle+c_{1}|1\rangle_{V}$, an external input radiation allows, after interaction, the transmission of an ideal coherent state $|\alpha\rangle$ inside a cavity mode for the excited state, the final “cat-like” state $|\psi\rangle_{1}=c_{0}|0\rangle|0\rangle+c_{1}|1\rangle|\alpha\rangle$ [7, 59, 60]. The resulting output signal has two possible unambiguous intensity outcomes 00 or $I=|\alpha|^{2}$ in absence of a superposition. Considering $N$ measurements without a detector reset in between two events, the average total intensity is $\overline{I}=N|c_{1}|^{2}I$ and the fluctuations $\overline{\delta^{2}I}=N[|c_{1}|^{2}(1-|c_{1}|^{2})]I^{2}$. Clearly, the relative fluctuations shrink to zero for large $N$. In this sense, $|c_{0}|^{2}$ and $|c_{1}|^{2}$ are the probability amplitude that is measurable after a counting over many events. A more thorough theoretical description of a qubit and resonator should have been used involving a quantum field approach using a output Fock basis as in Section III.2 [53].
@@ V.1 General considerations
In its orthodox formulation, quantum mechanics does not support the principles of locality and realism, and a wealth of experimental works inspired by Bell’s inequality test and conducted for over five decades [61, 45, 46, 47, 62, 63, 64, 65, 66, 67, 68, 69], led their authors to rule out both principles. Interestingly, the majority of these experiments involve the polarization direction of light as the physical quantity of interest, while Bell’s theoretical work considered 1/2 spins and a Stern-Gerlach-like experiment [70]. Further, and as important, the cited experiments have loopholes (such as, e.g., detection, locality, memory effects and statistics [71]) which preclude absolutely certain conclusions. In 2015, a work reported a loophole-free Bell inequality violation with an experiment involving electron spins [72]. The authors conclude that “[their] result places the strongest restrictions on local realistic theories of nature to date”; and this is indeed the case. But one may still argue that the principles of locality and realism are not disproved yet, and that other loopholes waiting to be identified may have been overlooked.
Here, for clarity, we neither claim that quantum correlations or entanglement at long distance are not possible, nor that Bell states do not exist. The singlet or triplet states in the helium atoms are examples of such entangled electron spin pairs. However, performing any Bell-like experiment in a framework based on a Fock space description imposed by quantum field theory is questionable once it is established that the collected data correspond to events that are not well separated because of their non-vanishing space and time correlations. Hence, there cannot be independent measurements in the context of a Bell experiment. This is a physics-based loophole, which never was examined. Such a type of experiment would certainly be feasible if there would be a means to distinguish identical particles, which would substantiate the assumption that two consecutive measurement events can be well separated. But given the experimental reality of non-discrimination between the detector and the setup, the current interpretation of Bell experiments does not involve the physical description of the detector and the source.
Without the Born’s postulate, the quantum description of the Bell experiment needs be reformulated. Particularly, the impossibility of a perfect interpretation of a noisy measurement prevents its feasibility, and hence any of two non-local correlated observed outcomes (spin up, spin down) is necessarily the result of many non-separable events. Under these conditions, locality, causality, and completeness are restored and there is no need of additional, hidden variables as shown below.
@@ V.2 Influence of the source thermal noise on photon pair production
A set of pair states produced in an experiment is quite different from what John Bell assumed in his work [73]. Using a continuous variable formalism associated to photons, we can factorize the polarized state symmetrically if the EPR source is not noisy. But in this case, all outcomes in both polarization directions would occur simultaneously contrary to observations where only one direction is observed. A possible explanation for such a discrimination is the inherent thermal noise of the source, like a beam of calcium atoms [45, 46, 47] or a nonlinear crystal [49] that generates additional classical photon flux fluctuations blurring the intrinsic quantum ones, creating a set of random pulse events. In a particle picture, an itinerant photon shows a Brownian-like dynamics as it interacts with atoms whose motion relative to the detector produces a Doppler shift with a relative frequency change $\sim\sqrt{1/\beta m}/c\sim 10^{-6}$ where $m$ is the mass of the atoms of the source, $c$ the speed of light, and $\beta$ the inverse thermal energy reflecting the temperature statistically determined from the average over the speeds of the atoms in motion.
@@
We can consider it as a stochastic variable whose relative average variation is given by the Doppler shift of a moving atom within the nonlinear crystal or the velocity distribution of calcium atoms. This results in a variance given by $\overline{\eta^{2}_{k,\epsilon,\epsilon^{\prime}}}(t)=\delta_{\epsilon,\epsilon^{\prime}}(\hbar\omega_{k})^{2}/\beta mc^{2}$. Comparing with the detuning and the nonlinear coefficient, these stochastic terms have an effect $(\omega_{k_{0}}/\Delta_{k_{0}})\sqrt{1/\beta mc^{2}}$ which can be of the order of unity or greater. As a consequence, such important contributions lead to potentially strong local variations in the intensity for creating histories of polarization mimicking hidden variables.
A basic formulation of the Bell experiment relies on the multi-mode squeezed state
@@
Assuming instead that pulses with photon numbers greater than unity are negligible with only one wavevector mode, we could employ the naive approximation of projecting the state onto the particle number pair of Eq. (20) to recover the symmetric Bell entangled state:
$|\psi\rangle\sim\frac{1}{\sqrt{2}}\left(\hat{a}^{\dagger}_{H}\hat{b}^{\dagger}_{H}+\hat{a}^{\dagger}_{V}\hat{b}^{\dagger}_{V}\right)|0\rangle$ (25)
where we use the simplification: $\hat{a}_{\epsilon}=\sum_{k}\hat{a}_{k,\epsilon}$ and $\hat{b}_{\epsilon}=\sum_{k}\hat{b}_{k,\epsilon}$. The projection postulate then allows different polarisation outcomes. The expression of the density matrix in Eq. (23) shows that using a QED formalism, genuine quantum correlations appear among different momentum modes but not among different localizations in space as Eq. (25) leads to think. Hence the question: if starting from a local quantum state that evolves according to quantum field theory, is it possible to produce a genuine entanglement at two distant points in space resulting in a non-local quantum state? According to [76], the existence of such non-local state implies non locality. Most of the textbooks assume their existence but the production of non locality starting from locality is never addressed on a firmly grounded basis.
@@ VI Conclusion
We also discussed with basic illustrative examples the use of the second quantization multi-particle formalism to avoid the pitfalls of the wavepacket collapse in the Schrodinger’s cat paradox. While the use of statistics is inherent to quantum physics experiments, this does not necessary entail that probability is a basic feature of quantum mechanics. It is inherent because the intrinsic uncertainty associated to common observables including, e.g., position, momentum, or particle number, requires a large number of data to be properly defined as a quasi-noiseless macroscopic measurement, which is essential for a conclusive interpretation. This amounts to adopting a thermodynamics formulation of the problem, where only macroscopic quantities are measurable since any single-event signal is too noisy to ensure a valuable characterization. For example, the observation of the electronic cloud in a solid crystal under a scanning tunnelling microscope is not a reality proposed by quantum mechanics but a representation of the total electronic density distribution resulting from of a macroscopic measurement current. From our proposed many-body modeling including the thermal noise of the source and of the detector, we conclude that the experiments in Refs. [45, 46, 47, 51, 49] are not proven either to correspond to the idealistic Bell’s proposal. Any genuine modeling of a quantum correlation experiment, with the hope of more deterministic predictions, should ideally include: 1) The full formalism of quantum electrodynamics with a many-body formalism highlighting indistinguishability; 2) A correct modelling of the source (e.g., a calcium beam with the two lasers); 3) An accurate quantum model of the detector which operates in a finite time and whose finite size imposes a lower bound over the counting rate preventing genuinely distinguishable (time-separated) single-event detection. A procedure along these steps results in a deterministic quantum description including locality and causality.
@@
More complex experiments that necessitate modeling beyond the traditional scattering approach using the S matrix, such as, e.g., the avalanche effect, need also new theoretical methods to address properly the exact quantum dynamics involved. For instance, the formalism of transmission within waveguides usually derived from classical electromagnetism has been extended using quantization [53]. These developments raise the fundamental questions on the possibility of creating a superposition of macroscopic “cat” states leading to large quantum fluctuations and therefore not measurable. In this work, we have shown the difficulty to produce such a “multiverse” state given the presently available technology; but a more accurate proof is required to establish this observation on a more formal basis. The non existence of such states amounts to conjecturing that any macroscopic observable has its set of possible eigenvalues restricted to regions within its fluctuations, discarding any extreme value far from its average. Such a statement should ideally be proven starting from a quantum field model.
@@
As a final remark regarding the discrimination in the theoretical description between the object of study and the measurement apparatus, it is worth recalling that when electric circuits were discovered in the time of Alessandro Volta in 19th century, both the circuit (e.g. voltaic pile + metal wires) and the measurement device (e.g. compass or galvanometer) had eventually required the electromagnetic theory for a full understanding of an experiment.